OpenStudy (anonymous):

S(t) = 68 − 20 log (t + 1), t ≥ 0 What was the average score after 4 months? after 24 months? After what time t was the average score 50%?

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

Round to the nearst two decimal places

OpenStudy (anonymous):

calculator exercise. replace t by 4, and 24. i will write it if you like

OpenStudy (anonymous):

is this log base ten?

OpenStudy (anonymous):

i dont get it

OpenStudy (anonymous):

t = time in months yes?

OpenStudy (anonymous):

so for the first one put \[S(4)=68-20log(4+1)=68-20log(5)\]

OpenStudy (anonymous):

in other words put 4 where you see a t and then compute

OpenStudy (anonymous):

you need a calculator to find log(5)

OpenStudy (anonymous):

i got 54.02

OpenStudy (anonymous):

let me try

OpenStudy (anonymous):

\[S(24)=68-20log(25)=...\]

OpenStudy (anonymous):

i let me know if you have an answer

OpenStudy (anonymous):

okay me to so for 24 months its 40.04

OpenStudy (anonymous):

sound right

OpenStudy (anonymous):

hold on i will compute

OpenStudy (anonymous):

yup that is what i got

OpenStudy (anonymous):

now we have to find when S = 50. do you know how to do that one?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

first replace S(t) by 50 \[50=68-20log(t+1)\] then get the log by itself on one side \[20log(t+1)=68-50=18\]

OpenStudy (anonymous):

ok so far?

OpenStudy (anonymous):

no , why did you put 68-50

OpenStudy (anonymous):

ok let me go step by step. \[50=68-20log(t+1)\] add \[20log(t+1)\] to both sides to get \[20log(t+1)+50=68\] then subtract 50 from both sides to get \[20log(t+1) = 68-50=18\] then divide by 20 to get \[log(t+1)=\frac{18}{20}=\frac{9}{10}=.9\]

OpenStudy (anonymous):

i did not need to add log(t+1) to both sides, but i prefer dealing with positive things when i am trying to solve. i could have subtracted 68 from both sides and gotten the same answer working with negative numbers. matter of preference.

OpenStudy (anonymous):

so we have arrived via some algebra at \[log(t+1)=.9\] is that ok so far?

OpenStudy (anonymous):

yes i think im getting it

OpenStudy (anonymous):

these were just algebra steps to get log(t+1) by itself. same steps i would have used to solve for x if the equation was \[50=68-20x\]

OpenStudy (anonymous):

you would get \[x=.9\] and here \[x=log(t+1)\] so \[log(t+1)=.9\]

OpenStudy (anonymous):

we are not done yet, though because we have to solve for t.

OpenStudy (anonymous):

okay im following you

OpenStudy (anonymous):

\[log(x)=y \] \[10^y=x\]

OpenStudy (anonymous):

what i meant to say was \[log(x)=y\] means the same thing as \[10^y=x \]

OpenStudy (anonymous):

here we have \[log(t+1)=.9\] so \[10^{.9}=t+1\]

OpenStudy (anonymous):

to find \[10^{.9}\] in need a calculator. then to get t by itself subtract 1.

OpenStudy (anonymous):

let me know what you get.

OpenStudy (anonymous):

6.94 sound right

OpenStudy (anonymous):

its wrong???